print this page
Print page
email to a friendEmail to friend

Daily Updates: January
23 24 25 26 27 28 29
30 31
Daily Updates: February
1 2 3 4 5
6 7 8 9 10 11 12

View Today's Slideshow!

brokenclouds weather

Broken Clouds
80°F (26.7°C)
Latitude: 9 deg 49’N
Longitude: 104 deg 17’W
Wind Direction: NE
Wind Speed: 14 Knots
Sea State: 3Swell(s) Height: 9 Feet
Sea Temperature: 81°F (27.2°C)
Barometric Pressure: 1012.0 MB
Visibility: 12 Nautical Miles

what's to eat today?

Scrambled eggs
Bacon, sausages
Banana muffins
Fresh fruit
Cream of Wheat

Battered fish
Buffalo wings
Macaroni and cheese
Chicken and vegetable soup
Salad bar
Seafood pasta

Burrito Buffet
    Chicken, shrimp, beef
    beans, rice, cheese,
    tomato, onion, salsa, olives,
    chili's, sour cream, lettuce, guacamole
Jell-O with fruit

Daily Update: Dive 3527
February 2, 2000
By Dr. Dan Fornari and Sam Dean

The morning dawned with gray puffy clouds lining the horizon and a sliver of new moon high in the sky. With Alvin’s penetrator repaired (see yesterday's dive for more on what happened) and all the pre-dive checks completed, Gary Comer and Bob James followed BLee Williams, the pilot, into Alvin so that they could dive to the hydrothermal vents near 9° 50’N Latitude. They will continue to collect data for several scientists who have programs that involve vent mapping and fluid sampling, as well as digital imaging of different volcanic and vent features in this area. BLee, Bob and Gary will also recover some temperature probes that have been measuring the temperature at the black smoker vents since May of 1999. The plan is to recover these probes and then redeploy them on the last dive of this cruise.

The Towed Camera Sled was hauled on board at 0430 hours this morning. Even after charging the batteries on the Sled for 11 hours, though, it still ran out of power after 7 hours. It seems that some more tweaking may be needed in order to draw more “juice” out of the batteries. This is only a minor setback. With the digital camera on the Sled returning between 1600-1800 crystal-clear pictures of the ocean bottom each night, Jenny, Greg, Margo and Del have plenty of images to compile and archive every day, including one photo of an animal with lots of legs that you may recognize!

And now for the ping pong tournament update! The entire first round has been played, and most of the second round as well. With strong play all-around, Captain George Silva and Paul Vinitsky have both moved into the semifinals. Stay tuned to find out who fills in the other two spots left in the semifinal bracket!

After the sub got back on deck at 1715 hours, and Bob and Gary were given a thorough hosing down for being new divers, the ship steamed down to 9° 37’N Latitude to continue the Towed Camera surveys. Tow #5 is planned for tonight.

Before putting the Camera Sled over the side, Dan Fornari tried to recover a sample of the lava on the seafloor using a spring loaded grab sampler. A picture of the grab is shown in today's slide show. The way it works is that the grab is cocked open, and then lowered over the side of the ship using the hydrographic winch which has a 5/16 inch steel wire that attaches to the grab. When the grab hits the ocean bottom it is designed to snap shut and “grab” a sample. Although normally used to pick up soft sediment, Dan was testing this device for use on an upcoming cruise (Cruise #3 in Dive and Discover!) to sample the rocks and biology on the East Pacific Rise crest. After its test run, though, the grab was brought on board and we found out that the jaws didn’t snap shut. In science, as in life, not everything works perfectly on its first try. Dan and the ship’s crew have some ideas on what happened, or rather didn’t happen, and we’ll see if they can work out the kinks and have better success with the grab in a few days.

Dive Summary
On Bottom: 0924 hours
Off Bottom: 1537 hours
Maximum Depth: 2508 meters

BLee, Bob, and Gary worked at three high temperature hydrothermal vents and one low temperature vent. Starting at the south end of the Bio-Geo Transect, a line of seafloor markers that extends for about 1500 meters in the axial trough, they went to “Tubeworm Pillar”, “Marker 141” (a low temperature vent), “Io”, “Ty”. At the vents they were able to see underwater communities teeming with life! They took some samples of hydrothermal vent fluids at the Tubeworm Pillar using gas-tight bottles for Dr. Marvin Lilley at the University of Washington and Prof. Karen Von Damm, a geochemist at the University of New Hampshire. Marv and Karen have been studying these vents since the 1991 eruption. They also recovered the temperature probe and a sulfide sample. When we got the probe back we noticed that there were baby Tevnia tubeworms growing on the handle of the probe!

One surprise in today’s dive was that one of the high temperature vents, Ty, had tumbled over. BLee was able to locate the temperature probe in the rubble. We will download the data tomorrow and find out when the chimney fell over.

In addition to taking spectacular digital photos and videos of the vents and communities of animals around them, they were able to recover some great samples from the area. They were able to collect some samples of the giant Riftia tubeworms, mussels, and crabs. They also collected a couple more lava samples, including a lava pillar just north of the Marker 141 area from a region of sheet flow lava adjacent to the west wall of the axial trough. All in all a terrific dive and a great collection of samples!